OpenStudy (anonymous):

x^2+8x-2=0 solve using quadratic formula

OpenStudy (anonymous):

x=(-8+-sqrt(64-4*(-2))/2=(-8+-sqrt(72))/2; x=-4+-sqrt(72)/2; x=-4+-sqrt(9*8)/2; x=-4+-3sqrt(2) so x=-4+3sqrt(2) and x=-4-3sqrt(2)

OpenStudy (anonymous):

\[\text{Quadratic Formula: x} = \frac {-b ± \sqrt{b^{2} - 4ac}}{2a}\]

OpenStudy (anonymous):

\[x = \frac{-8 ± \sqrt{8^{2} - 4(1)(-2)}}{2(1)}\]\[x = \frac{-8 ± \sqrt{64 + 8}}{2}\]\[x = \frac{-8 ± \sqrt{72}}{2}\]\[x = \frac{-8}{2} ± \frac{6\sqrt{2}}{2}\]\[x = -4 ± 3\sqrt{2}\]

OpenStudy (anonymous):

thanks!